# Deep Reinforcement Learning-Based Active Flow Control of an Elliptical Cylinder: Transitioning from an Elliptical Cylinder to a Circular Cylinder and a Flat Plate

### v1

@@ III.2 Effectiveness of AFC
This section describes the performance of AFC using the DRL algorithm, which manipulates the magnitude of synthetic jets to achieve targeted flow control. Within a flow configuration at $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, as well as to cylindrical and flat plate configurations. The evaluation of the control’s effectiveness and performance is based on several criteria: comparisons of the velocity fields before and after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
Figure 7: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline simulation. Left side of each row: result with active flow control. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$.
Fig. 7 and Fig. 8 provide flow field snapshots of elliptical cylinders under various $A$, encompassing both baseline (uncontrolled) and controlled scenarios. This section primarily describes the flow structures observed in the baseline cases. For elliptical cylinders with $Ar=2$, the baseline flow field inherently displays high stability, with no vortex shedding observed. Consequently, the AFC did not produce a noticeable effect. As the aspect ratio decreases to $Ar=1.75$, vortices begin to detach alternately from both sides of the elliptical cylinder at reduced frequencies and intensities. These detaching vortices organize into a relatively orderly and stable vortex street. By applying AFC based on DRL, the vortex shedding in the wake of the elliptical cylinder is completely suppressed, resulting in a stabilized wake field. At an aspect ratio of $Ar=1.5$, an alternating vortex structure forms behind the elliptical cylinder. These structures possess certain strengths and frequencies, yet the extent of vortex shedding remains comparatively mild. Further reduction in the aspect ratio to $Ar=1.25$ leads to a more pronounced and frequent shedding of vortical structures from the wake, indicating a transition towards more significant flow instability. The recirculation bubble downstream of the elliptical cylinder diminishes, oscillates, and becomes unstable. When the aspect ratio is $Ar=1.5$ and $Ar=1.25$, a stable, regular alternating vortex structure forms behind the elliptical cylinder. After the application of AFC, the shedding of vortices behind the elliptical cylinder is completely suppressed, transitioning the wake field into a stable flow structure. At $Ar=1$ and $Ar=0.75$, the area of the recirculation bubbles formed behind the elliptical cylinder decreases, leading to more prominent and frequent vortex shedding within the wake, indicating a transition towards significant flow instability. Following the implementation of AFC, the alternating vortex street behind the elliptical cylinder is transformed into a stable wake field.
Fig. 7 depicts the instantaneous velocity fields both baseline and post-implementation of control measures across a range of $Ar$ from 2 to 0.75 for elliptical cylinders. At $Ar=2$, the baseline wake field is notably stable. However, when the $Ar$ is reduced below 1.75, alterations in structural parameters provoke enhanced vortex shedding behind the ellipse, characterized by escalating frequency and intensity as the $Ar$ diminishes. The efficacy of AFC is underscored through the deployment of a control strategy, which effectively attenuates vortex shedding behind the elliptical cylinder for $Ar=1.75$, 1.5, 1.25, 1, and 0.75. This intervention demonstrates the significant capabilities of the DRL-based AFC technology in manipulating vortex shedding within the wake fields of elliptical cylinders across varying $Ar$, thereby enhancing the flow stability of the configurations under study.
Figure 8: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline flow. Left side of each row: result with active flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
As the $Ar$ is further reduced to 0.5, 0.25, and 0.1, eventually transitioning to a flat plate configuration, the vortex shedding phenomena behind the elliptical cylinder become more frequent and intense. It is evident that decreasing the $Ar$ leads to a significant increase in the circulation entering the wake, as the flow velocity increases when water moves around the top of the cylinder. This observation aligns with the findings of Thompson et al.Thompson et al. (2014). This underscores that as the $Ar$ of an elliptical cylinder gradually diminishes towards zero, it induces an enhancement in the flow instability of the wake field, consequently leading to the release of more energy in the form of vortices. This dynamic highlights the critical role of structural parameters in influencing the energy distribution and stability of the flow field as it interacts with varying geometrical configurations.
However, when utilizing the AFC technique to manipulate the flow field, the control outcomes differ from those observed at $Ar=0.75$ and higher $Ar$; vortex shedding behind the elliptical cylinder is not completely suppressed. At $Ar=0.5$, analysis of the instantaneous velocity field reveals that the recirculation area behind the elliptical cylinder is significantly elongated, and the shedding of vortices is notably subdued. Nevertheless, the flow field has not fully stabilized, as there remains a continual detachment of alternating vortex structures from behind the elliptical cylinder, resembling elongated vortex formations. At $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is similarly elongated as seen at $Ar=0.5$, but with a notable difference: the bubble now exhibits significant oscillations. In the controlled wake field, the phenomenon of flow separation following fluid passage around the elliptical cylinder and subsequent reattachment is also observable. Despite these dynamics, the instability of the downstream velocity field has been relatively stabilized compared to the baseline state, indicating the effectiveness in delaying vortex shedding. At $Ar=0.1$ or 0, compared to the baseline condition, the control of the flow field has primarily delayed rather than completely suppressed vortex shedding behind the elliptical cylinder. The recirculation bubble behind the cylinder is elongated, yet it exhibits significant oscillations. This indicates that while the AFC has extended the duration before vortex detachment, it has not fully stabilized the vortex structures within the wake.
In summary, at higher $Ar$ ($Ar=0.75$, 1, 1.25, 1.5, 1.75), the DRL-based control strategy can fully suppress vortex formation, thereby stabilizing the wake field. However, at lower $Ar$ ($Ar=0.5$, 0.25, 0.1, 0), the effectiveness of the control strategy is limited to mitigating vortex shedding within the wake field to some extent. It delays the vortex detachment process and enlarges the recirculation area behind the elliptical cylinder, but oscillations in the recirculation bubble persist, and the downstream flow field continues to exhibit signs of instability.
Figure 9: Comparison of $C_{D}$ and $C_{L}$ before and after control.(a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
@@
Fig. 9 presents a comparative analysis of the $C_{D}$ and $C_{L}$ before and after control intervention over $Ar$ ranging from 2 to 0.75. At an $Ar$ of 2, both $C_{D}$ and $C_{L}$ are stable, indicating that AFC does not have any effect. However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, upon activating synthetic jets control, there is an immediate, sharp response in $C_{D}$, which quickly descends to a value near its minimum. After half the total control duration, $C_{D}$ stabilizes and remains constant at this minimal value. Similarly, $C_{L}$ experiences significant fluctuations at the start of control but stabilizes near zero after half the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ are detailed in Tab. 2. These observations substantiate the effectiveness of the active flow control strategy underpinned by DRL algorithms. The activation of control swiftly alters the periodic behavior of both $C_{D}$ and $C_{L}$, facilitating the stabilization of $C_{D}$ and $C_{L}$ in the subsequent stages. This demonstrates the capability of the DRL-based active flow control approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
@@
Fig. 10 illustrates the $C_{D}$ and $C_{L}$ for both baseline and controlled flow conditions around an elliptical cylinder at $Ar$ of 0.5, 0.25, 0.1, and 0. Under AFC, at $Ar=0.5$ and $Ar=0.25$, both $C_{D}$ and $C_{L}$ are reduced compared to the baseline flow, yet they do not achieve the stable state observed at $Ar=0.75$ and higher. Initially, both the $C_{D}$ and the $C_{L}$ experienced abrupt changes. However, they did not fully stabilize in the subsequent period. In the latter half of the control duration, $C_{D}$ and the $C_{L}$ exhibited fluctuations, but both the amplitude and frequency of these fluctuations were lower than those observed in the baseline flow. At $Ar=0.1$ and $Ar=0$, the controlled flow experiences a sharp decrease in the drag coefficient at the onset of control, followed by oscillations that persist at a lower frequency and amplitude than the baseline flow. Although complete stability is not achieved, the average value of the $C_{D}$ is significantly reduced. The $C_{L}$ undergoes an initial sharp decrease and approaches near zero, and after half the total duration of control, it oscillates at a lower frequency.
Figure 10: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline flow. Left side of each row: result with active flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
Figure 11: Action, which is the mass flow rate of the synthetic jets. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
@@
Fig. 11 illustrates the time history curve of the action (the mass flow rate of synthetic jets), When performing AFC using DRL for flow around an elliptical cylinder with $Ar$ between 0 and 2. When $Ar=2$, the AFC is ineffective, and the value of the action, pertaining to the control input, remains close to zero. From $Ar=1.75$ to $Ar=0.75$, during the initial phases of control, the agent selects larger action values, corresponding to higher mass flow rates of synthetic jets. This indicates a deliberate strategy to assert flow control by introducing significant external energy inputs. As control progresses, the agent gradually diminishes the action values, thereby reducing the mass flow rate of the synthetic jets. Subsequently, in the mid to later stages, the agent’s action values converge towards zero, corresponding to a near-zero mass flow rate of synthetic jets. This behavior signifies the agent’s optimization towards an energy-saving state, maintaining flow characteristics with minimal control input. When $Ar=0.5$ and $Ar=0.25$, the value of the action initially surges but does not stabilize near zero in the later stages. For $Ar=0.1$ and $Ar=0$, the action value remains in a state of periodic fluctuation.
@@
Overall, the results presented in Fig. 7 and Fig. 9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=0.75$ to $Ar=1.75$. By progressively decreasing the mass flow rate of synthetic jets, the strategy achieves flow control, reduces drag and lift forces, and does so with reduced external energy input. However, at $Ar=0.5$ and $Ar=0.25$, the energy efficiency of the AFC strategy diminishes. At $Ar=0.1$ and $Ar=0$, the external energy utilized by the control strategy is relatively high compared to the inlet velocity, which will be detailed in Tab. 2.
@@
Figure 12: The learning curve of PPO algorithm training when $Ar=2$ to 0. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
@@
In Fig. 12, we observe the learning curve during training of flow around an elliptical cylinder with different $Ar$. At $Ar=2$, the reward function value never reaches a positive figure, indicating that the DRL training did not achieve its objectives. For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4000 episodes converges near the maximum reward value. At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3000 episodes and eventually settles into a stable plateau. For $Ar=1$ and $Ar=0.75$, there are significant initial fluctuations with the learning curve rising swiftly, indicating rapid progress in policy learning. After 2000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2000 and 5000 episodes. This shows that the PPO algorithm has successfully learned an effective strategy, exhibiting commendable convergence and stability in performance. At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4000 episodes, with the reward values stabilizing between 4000 and 10000 episodes. At $Ar=0.25$, training approaches convergence after 8000 episodes, reaching a total of 20000 episodes, with a mid-to-late plateau marking stable convergence. At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12000 episodes.
@@
Overall, the DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process. Training also achieves stable convergence within 5000 episodes for $Ar=1.5$ and $Ar=1.75$. However, for $Ar=0.5$, 0.25, 0.1 and 0, as the flow phenomena become increasingly complex, the control challenge intensifies, leading to a noticeable slowdown in convergence speed. This indicates that for flow scenarios with higher degrees of instability, the convergence performance of DRL training is compromised, requiring a larger number of training iterations to reach stable convergence.
Tab. 2 presents the effectiveness of active flow control techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0.1 to 2. It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change. Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate. Initially, at $Ar=2$, the wake field stability behind the elliptical cylinder is observed to be relatively high, resulting in no significant alteration in the drag experienced by the elliptical cylinder. When $Ar$ is reduced to 1.75, there is a noticeable decrease in drag around the elliptical cylinder by 0.9%. With further reduction in $Ar$, the rate of drag reduction progressively increases. In their study, Protas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations. This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control. In this research, the flow field is controlled by utilizing "external excitation," such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance. As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag.
@@
In Tab. 2, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case. These results highlight the efficacy of the DRL-based AFC technology in effectively suppressing both $C_{D}$ and $C_{L}$ within the $Ar$ range of 1.75 to 0.75, demonstrating the efficiency of the DRL-based intelligent control strategy. However, as $Ar$ decreases to 0.5, 0.25, 0.1, and 0, even though the $C_{D}$ reduction rates significantly increase to 26.9%, 35.9%, 37.6%, and 43.6%, the $C_{L}$ reduction rates are 68.0%, 66.1%, 50.2%, and 60.0% respectively. This indicates that while the DRL-based control strategy achieves significant reductions in $C_{D}$, its effect on suppressing $C_{L}$ is limited and does not completely eliminate $C_{L}$ around the elliptical cylinder. As the $Ar$ further decreases, the AFC technique based on the DRL algorithm does not fully suppress vortex shedding behind the elliptical cylinder, but merely delays the shedding process. The reduction in $C_{L}$ coefficients aligns with the flow snapshots presented in Fig. 8, showing that vortex shedding still occurs, albeit with controlled frequency and intensity. These phenomena suggest that at very low $Ar$ values, the control strategy faces challenges in managing more complex and unstable flow conditions.
@@
The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers. Typically, approximately 93% of the drag induced by the shedding of vortices is suppressed by the discovered control laws, with minimal jet intensities required for drag reduction. For elliptical cylinders with aspect ratios $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.

### v2

@@ III.2 Effectiveness of AFC
This section describes the performance of AFC using the DRL algorithm, which manipulates the magnitude of synthetic jets to achieve targeted flow control. At $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, from elongated body to circular cylinder and then to a normal flat plate. The evaluation of the control’s effectiveness and performance is based on several criteria: intensity of vortex shedding after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process. Fig. 7 and Fig. 8 display the velocity magnitude snapshots of elliptical cylinders with $Ar$ ranging from 2 to 0 under baseline and controlled conditions. More specifically, Fig. 7 shows those for $Ar$ from 2 to 0.75, the vortices in the wake field of the controlled elliptical cylinders can be completely suppressed. In contrast, Fig. 8 illustrates those for $Ar$ from 0.5 to 0, the vortices in the wake field of the controlled elliptical cylinders are mitigated by not fully suppressed.
Figure 7: Comparison of the velocity magnitude snapshots for an elliptical cylinder before and after control. Each row displays: on the left, the baseline simulation; on the right, the AFC results. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$.
Starting from Fig. 7(a) for the elliptical cylinder with $Ar=2$, the baseline flow field (left) exhibits high stability with no observable vortex shedding. We applied a DRL-based flow control to this stable, non-shedding wake field, and the controlled flow field (right) behind the elliptical cylinder remained stable and nearly identical to the baseline. This sanity test indicates that the DRL-based model is consistent with flow physics and does not generate any unreasonable perturbation during the control process. As the $Ar$ decreases to $Ar=1.75$, the application of DRL-based AFC technology successfully suppresses the vortices that originally separated alternately from both sides of the elliptical cylinder. Further reducing the $Ar$ results in a more pronounced and frequent shedding of vortex structures in the uncontrolled wake, while the recirculation bubble downstream of the elliptical cylinder gradually decreases and becomes oscillatory. Although the flow instability gradually increases, under the conditions of $Ar=1.5$, 1.25, 1, and 0.75, the alternating vortex streets behind the elliptical cylinders are transformed into a stable wake field after the implementation of AFC in all cases. The DRL-based AFC technology demonstrates significant vortex shedding suppression capabilities under these conditions, highlighting the effectiveness of the AFC control strategy in enhancing the flow stability of the wake field.
Figure 8: Comparison of the velocity magnitude snapshots for an elliptical cylinder before and after control. Each row displays: on the left, the baseline simulation; on the right, the AFC results. (a) $Ar=0.5$; (b) $Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
In Fig. 8, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the baseline flow becomes increasingly unstable, and the area of the recirculation region gradually decreases while the strength and frequency of vortex shedding increase. When $Ar=0.5$, the frequent alternating vortex shedding is effectively suppressed after flow control is applied. The recirculation region behind the elliptical cylinder is significantly elongated, and the shedding of vortices is controlled, although the wake flow still exhibits alternating oscillations. For the controlled flow at $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is elongated, resembling the case at $Ar=0.5$. However, there is a notable difference: the recirculation bubble now demonstrates prominent oscillations. The vortex shedding behind the cylinder continues alternately, akin to the elongated vortex structures being stretched longitudinally. Despite these instabilities, vortex shedding in the wake fields at $Ar=0.5$ and $Ar=0.25$ is reasonably mitigated compared to the baseline state. At $Ar=0.1$ or 0, compared to baseline conditions, the recirculation bubbles behind the cylinder are elongated, and the initiation point of vortex shedding is extended further downstream. We do not observe complete suppression of vortex shedding due to the highly chaotic nature of these wake behind vertical thin bodies. Nevertheless, the vortex shedding is reduced to a lower frequency and intensity. These observations reflect the effectiveness of the control strategies in suppressing vortex shedding and stabilizing the wake field.
Figure 9: Comparison of $C_{D}$ and $C_{L}$ before and after control. The background color represents the unstable stage of control, and the absence of background color means that the corresponding drag or lift has been in a stable stage during this period. (a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
@@
Fig. 9 presents a comparative analysis of the $C_{D}$ and $C_{L}$ before and after control intervention over $Ar$ ranging from 2 to 0.75. At an $Ar$ of 2, both $C_{D}$ and $C_{L}$ are stable, indicating that AFC does not have any effect. However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, upon activating synthetic jets control, there is an immediate, sharp response in $C_{D}$, which quickly descends to a value near its minimum. After half the total control duration, $C_{D}$ stabilizes and remains constant at this minimal value. Similarly, $C_{L}$ experiences significant fluctuations at the start of control but stabilizes near zero after half the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ are detailed in Tab. 2. These observations substantiate the effectiveness of the AFC strategy underpinned by DRL algorithms. The activation of control swiftly alters the periodic behavior of both $C_{D}$ and $C_{L}$, facilitating the stabilization of $C_{D}$ and $C_{L}$ in the subsequent stages. This demonstrates the capability of the DRL-based AFC approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
@@
Fig. 10 illustrates the $C_{D}$ and $C_{L}$ for both baseline and controlled flow conditions around an elliptical cylinder at $Ar$ of 0.5, 0.25, 0.1, and 0. Under AFC, at $Ar=0.5$ and $Ar=0.25$, both $C_{D}$ and $C_{L}$ are reduced compared to the baseline flow, yet they do not achieve the stable state observed at $Ar=0.75$ and higher. Initially, both the $C_{D}$ and the $C_{L}$ experienced abrupt changes. However, they did not fully stabilize in the subsequent period. In the latter half of the control duration, $C_{D}$ and the $C_{L}$ exhibited fluctuations, but both the amplitude and frequency of these fluctuations were lower than those observed in the baseline flow. At $Ar=0.1$ and $Ar=0$, the controlled flow experiences a sharp decrease in the drag coefficient at the onset of control, followed by oscillations that persist at a lower frequency and amplitude than the baseline flow. Although complete stability is not achieved, the average value of the $C_{D}$ is significantly reduced. Similarly, the $C_{L}$ undergoes an initial sharp decrease and approaches near zero, then later starts to oscillates, but at a lower frequency than the baseline $C_{L}$ oscillation.
Figure 10: Comparison of $C_{D}$ and $C_{L}$ before and after control. The background color indicates that the drag or lift coefficient is still in an unstable stage after flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
Figure 11: Action, which is the mass flow rate of the synthetic jets. The background color represents that the value of the action is in an unstable stage, and the absence of a background color means that the action has been in a stable stage during this period. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
@@
Fig. 11 illustrates the time history curve of the action (the mass flow rate of synthetic jets), When performing AFC using DRL for flow around an elliptical cylinder with $Ar$ between 0 and 2. When $Ar=2$, the AFC is negligible, and the value of the action, pertaining to the control input, remains close to zero. This is a strong indication that the DRL-based model is not generating nonphysical perturbations for a stable flow field. From $Ar=1.75$ to $Ar=0.75$, during the initial phases of control, the agent selects larger action values, corresponding to higher mass flow rates of synthetic jets. This indicates a deliberate strategy to assert flow control by introducing significant external energy inputs. As control progresses, the agent gradually diminishes the action values, thereby reducing the mass flow rate of the synthetic jets. Subsequently, in the mid to later stages, the agent’s action values converge towards zero, corresponding to a near-zero mass flow rate of synthetic jets. This behavior signifies the agent’s optimization towards an energy-saving state, maintaining flow characteristics with minimal control input. When $Ar=0.5$ and $Ar=0.25$, the value of the action initially surges but does not stabilize near zero in the later stages. For $Ar=0.1$ and $Ar=0$, the action value remains in a state of periodic fluctuation.
@@
Overall, the results presented in Fig. 7 and Fig. 9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=2$ to $Ar=0.75$. By progressively decreasing the mass flow rate of synthetic jets, the strategy achieves flow control, reduces drag and lift forces, and does so with reduced external energy input. However, at $Ar=0.5$ and $Ar=0.25$, the energy efficiency of the AFC strategy diminishes. At $Ar=0.1$ and $Ar=0$, the external energy utilized by the control strategy is relatively high compared to the inlet velocity, which will be detailed in Tab. 2.
@@
Figure 12: When $Ar$ of the elliptical cylinder is in the range of 2-0, the learning curve of PPO algorithm training. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
@@
In Fig. 12, we observe the learning curve during training of flow around elliptical cylinders with different $Ar$. At $Ar=2$, the learning curve has entered the plateau phase after the initial improvement phase, indicating that the DRL training has converged. For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4,000 episodes converges near the maximum reward value. At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3,000 episodes and eventually settles into a stable plateau. For $Ar=1$ and $Ar=0.75$, there are significant initial fluctuations with the learning curve rising swiftly, indicating rapid progress in policy learning. After 2,000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2,000 and 5,000 episodes. This shows that the PPO algorithm has successfully learned an effective strategy, exhibiting commendable convergence and stability in performance. At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4,000 episodes, with the reward values stabilizing between 4,000 and 10,000 episodes. At $Ar=0.25$, training approaches convergence after 8,000 episodes, reaching a total of 20,000 episodes, with a mid-to-late plateau marking stable convergence. At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12,000 episodes.
@@
The DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process. Training also achieves stable convergence within 5,000 episodes for $Ar=1.5$ and $Ar=1.75$. However, for $Ar=0.5$, 0.25, 0.1 and 0, as the flow phenomena become increasingly complex, the control challenge intensifies, leading to a noticeable slowdown in convergence speed. This indicates that for flow scenarios with higher degrees of instability, the convergence performance of DRL training is compromised, requiring a larger number of training iterations to reach stable convergence.
@@
Tab. 2 presents the effectiveness of AFC techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0 to 2. It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change. Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate. When the $Ar$ is set to 2, the baseline flow has reached a steady state. During the application of flow control, the agent does not use external energy to disrupt this stable flow field, which demonstrates the intelligence of the DRL-based AFC. This reflects the adaptability and intelligent decision-making of AFC control in maintaining flow stability. For elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case. These results highlight the efficacy of the DRL-based AFC technology in effectively suppressing both $C_{D}$ and $C_{L}$ within the $Ar$ range of 1.75 to 0.75, demonstrating the efficiency of the DRL-based intelligent control strategy. However, as $Ar$ decreases to 0.5, 0.25, 0.1, and 0, even though the $C_{D}$ reduction rates significantly increase to 26.9%, 35.9%, 37.6%, and 43.6%, the $C_{L}$ reduction rates are 68.0%, 66.1%, 50.2%, and 60.0% respectively. This indicates that while the DRL-based control strategy achieves significant reductions in $C_{D}$, its effect on suppressing $C_{L}$ is limited and does not completely eliminate $C_{L}$ around the elliptical cylinder. As the $Ar$ further decreases, the AFC technique based on the DRL algorithm does not fully suppress vortex shedding behind the elliptical cylinder, but merely delays the shedding process. The reduction in $C_{L}$ coefficients aligns with the flow snapshots presented in Fig. 8, showing that vortex shedding still occurs, albeit with controlled frequency and intensity. These phenomena suggest that at very low $Ar$ values, the control strategy faces challenges in managing more complex and unstable flow conditions.
@@
In their research, Protas and WesfreidProtas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations. This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control. In our research, the flow field is controlled by utilizing external excitation, such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance. As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag. This is the reason why the drag reduction rate gradually increases as $Ar$ decreases in Tab. 2.
The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers. Typically, approximately 93% of the drag induced by the shedding of vortices is suppressed by the discovered control laws, with minimal jet intensities required for drag reduction. In our research, the control strategy derived from the DRL-based AFC system dynamically responds in real time to variations in the flow state. This emphasizes its capability to adaptively adjust to changing conditions, demonstrating the robustness and responsiveness of the AFC framework. Furthermore, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.